Circular No. 8132

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4745 SAGITTARII N. M. Ashok and D. P. K. Banerjee, Physical Research Laboratory

(PRL), Ahmedabad, report near-infrared spectroscopy (range 1.08–2.30 μm) of V4745 Sgr with the Mt. Abu 1.2-m telescope (+ PRL Near Infrared Imager Spectrometer) on Apr. 28.9 UT. The spectra show strong H I emission lines at Paβ, Brγ, and Brackett 11 to 17. An additional weak emission line is seen at 1.316 μm, attributable to O I.

T. Kato, Kyoto University, reports that a spectrum of V4745 Sgr, taken by M. Fujii (Fujii Bisei Observatory) on May 8.79 UT, shows it to be a nova of Fe II class, with strong Balmer lines having P-Cyg profile.

V4745 Sgr brightened to visual mag ∼ 8.0 around May 8–12 before fading again. Photoelectric photometry of V4745 Sgr by A. C. Gilmore and P. M. Kilmartin at the University of Canterbury’s Mt. John Observatory (cf. IAUC 8127): May 4.610 UT, V = 9.51, U −B = −0.95, B −V = −0.03, V − R = +0.49, V − I = +0.23, air mass 1.10; 5.446, 9.50, −0.92, −0.03, +0.49, +0.22, 1.68; 11.600, 8.11, −0.75, +0.14, +0.21, +0.39, 1.09; 14.694, 9.22, −0.95, +0.01, +0.43, +0.31, 1.03; 15.588, 9.69, −0.97, +0.00, +0.61, +0.32, 1.09; 16.585, 10.34, −1.06, +0.03, +0.90, +0.44, 1.09. Additional CCD photometry: May 10.152, B = 8.37 (W. Liller, Viña del Mar, Chile); 10.155, V = 8.09 (Liller); 12.423, V = 8.35 (J. D. West, Mulvane, KS).

L. Kiss and A. Jacob, University of Sydney, report: “Optical spectra (0.05 nm/pixel, ranges 350–450 and 610–700 nm) were obtained during the rebrightening of this nova on May 12.66 and 13.67 UT with the Siding Spring 2.3-m telescope (+ DBS CCD spectrograph). The Balmer hydrogen series of emission lines shows complex multiple P-Cyg profiles. The Hα emission component has a 1600 km/s FWHM, while four distinct absorption components can be identified at −900, −1300, −1600, and −1800 km/s. The Hγ, Hδ, and Hϵ lines have similar profiles, with comparably strong Fe II emission lines at 417.6 and 423.3 nm, also with P-Cyg profiles.”

SUPERNOVAE 2003cg, 2003du, AND 2003ef Further to IAUC 8117, W. M. Wood-Vasey reports the following pre-

discovery magnitudes of SN 2003ef (cf. IAUC 8131) from NEAT Haleakala images: Feb. 16.47 UT, [19.6; Apr. 20.31, 19.5; 27.38, 17.5.

CCD R magnitudes by K. Hornoch, Lelekovice, Czech Rep.: SN 2003cg in NGC 3169, Mar. 23.994 UT, 14.5; 29.974, 13.9; Apr. 4.958, 14.3; 8.937, 14.3; 21.900, 14.8; 24.896, 14.9; May 3.883, 15.4. SN 2003du in UGC 9391, Apr. 24.937, 14.8; 27.910, 14.1; May 3.899, 13.5; 6.858, 13.4; 7.875, 13.4.

2003 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green